Suppose one basket has five white balls and seven black balls. A second basket has three...

  1. Suppose one basket has five white balls and seven black balls. A second basket has three white balls and twelve black balls. We will flip a fair coin and if it lands “heads” a ball from basket 1 is selected, and if it lands “tails” a ball from basket 2 is selected. Suppose that we have done this and that a white ball was selected, what is the probability that the coin landed tails?
Answer #1

probability of selecting white ball from basket 1 = 5+712

probability of selecting white ball from basket 2 = 3+12 5

probability that the coin landed tails =  12 3 125 60

Ans = 12/37

PLEASE LIKE IT RAISE YOUR THUMBS UP

IF YOU ARE HAVING ANY DOUBT FEEL FREE TO ASK IN COMMENT SECTION